Question text
A cannonball with a mass of 12.2 kg is fired from a cannon that has a mass of 511.1 kg. The cannonball is fired with a speed of 173.9 m/s, and it takes 0.009 seconds for the cannonball to leave the cannon.



Calculate the net force acting on the cannon.
1
Expert's answer
2020-04-15T10:16:48-0400

According to Newton’s third law, action is equal to counteraction. Therefore, with what force the cannon exerts on the cannonball with the same force (only in opposite direction), the cannonball exerts on the cannon. The cannonball, due to its accelerating force, acquires a momentum

(1) "\Delta P=m\cdot V=12.2 kg\cdot 173.9 m/s=2121.58 Ns" .

The average force that was needed for this is equal to

(2) "F=\frac{\Delta P}{\Delta t}=\frac{2121.58 Ns}{0.009 s}=2.36\cdot 10^5 N"

Answer: The net force acting on the cannon is "2.36\cdot 10^5 N"